# ORACLE: Explaining Feature Interactions in Neural Networks with ANOVA

### v2

@@ Abstract
We introduce ORACLE, a framework that explains neural networks on tabular and scientific design data. It fits ANOVA-style main and pairwise interaction effects to a model’s prediction surface. ORACLE treats a trained network as a black-box response, learns an orthogonal factorial surrogate on a discretized input grid, and uses simple centering and $\mu$-rebalancing steps to obtain main- and interaction-effect tables that remain $L^{2}$-consistent with the original model. The resulting grid-based interaction maps are easy to visualize, comparable across backbones, and directly connected to classical design-of-experiments analyses. On synthetic factorial and low- to medium-dimensional tabular regression benchmarks, ORACLE more accurately recovers ground-truth ANOVA interactions and hotspot structure than Monte Carlo SHAP-family interaction methods, as measured by ranking, localization, and cross-backbone stability metrics. In latent image and text settings, ORACLE instead delineates its natural scope, and our results indicate that grid-based ANOVA surrogates are most effective when features admit interpretable factorial structure, making ORACLE particularly well-suited to scientific and engineering tabular workflows that require stable, DoE-style interaction summaries.
@@ 1 Introduction
Modern neural networks often rely on rich interactions between features rather than purely additive structure, which makes them both powerful and difficult to interpret. Beyond marginal feature importance, practitioners increasingly need tools that describe how features act jointly, how such effects vary across the input space, and whether these patterns remain stable when the backbone model is changed. This need is particularly acute in scientific and engineering tabular settings, where inputs represent physical or design variables and stakeholders value stable, low-order interaction summaries that connect back to classical DoE practice. In this regime, analysis-of-variance (ANOVA) provides a natural formalism: it decomposes a response into main and interaction effects and, in classical design-of-experiments (DoE), uses factorial designs to estimate and visualize interactions with orthogonality and simple diagnostics.
In the neural network setting, most interaction explanations have grown out of the SHAP literature rather than DoE, relying on Monte Carlo sampling over feature subsets and often lacking a clear reference notion of an interaction surface. SHAP interaction indices and related methods can be computationally expensive, sensitive to sampling choices, and difficult to validate in terms of interaction detection, hotspot localization, and transfer across backbones. We instead revisit the DoE perspective and treat a trained neural network as a black-box response surface on which a factorial ANOVA surrogate is fitted. Our framework, ORACLE (Orthogonal and Residual ANOVA for Concordant $L^{2}$-Effect Explanations), learns an orthogonal surrogate on a discretized input grid and applies residual centering and $\mu$-rebalancing to recover main-effect and pairwise-interaction tables that remain faithful to the model in an $L^{2}$ sense. The resulting low-dimensional ANOVA interaction maps are directly visualizable, comparable across backbones, and aligned with DoE-style reasoning for tabular and scientific-design applications where features admit interpretable factorial structure and stakeholders value stable, low-order interaction summaries.
@@
    Formulate an ANOVA-based surrogate for neural predictors that yields orthogonal main and pairwise interaction tables with explicit $L^{2}$ semantics and associated interaction maps.
@@
    Define target ANOVA interaction strengths and maps from non-additive residuals on two-dimensional grids, providing a model-centric reference surface for evaluating interaction explanations.
@@
    Empirically compare ORACLE to SHAP-family interaction methods on synthetic factorial and real tabular regression benchmarks, highlighting regimes in which ANOVA surrogates yield more accurate and stable interaction explanations for low- to medium-dimensional tabular and scientific-design data.
@@ 2 Related Work
Recent work on global interaction explanations has largely built on the Shapley framework. Faith-Shap introduces an axiomatic Shapley interaction index as a canonical notion of global interaction strength (tsai2023faith), while SHAP-IQ, SVARM-IQ, and KernelSHAP-IQ provide Monte Carlo and kernel-based schemes to efficiently approximate any-order Shapley interactions in practice (fumagalli2023shap; kolpaczki2024svarm; fumagalli2024kernelshap). These algorithms and variants are consolidated and benchmarked in the shapiq library, which compares Shapley-based interaction estimators across models and datasets (muschalik2024shapiq). Complementary work analyzes how global feature effects decompose into main and interaction components (herbinger2024decomposing) and uses Shapley interactions to probe structured phenomena such as syntax and speech representations in large models (singhvi2024knowing). However, these approaches mostly return scalar interaction indices rather than explicit ANOVA interaction maps, and are rarely evaluated against an interaction oracle or for stability across backbones.
ANOVA and design-of-experiments views on interactions.
Functional ANOVA has re-emerged as a central tool for interpretable machine learning, motivating intrinsically interpretable models that decompose predictions into main effects and low-order interactions, as well as neural architectures that directly parameterize ANOVA components or screen interactions before fitting an ANOVA-style surrogate, enabling higher-order terms with statistical guarantees (hu2025interpretable; kohler2025achieving; choi2025meta; park2025tensor). In engineering design and global sensitivity analysis, interpretable neural networks are combined with functional ANOVA decompositions and grouped sensitivity indices to study complex physical interactions from observational data (dolar2024data). Complementary work in design-of-experiments couples factorial and fractional-factorial designs with modern machine learning to estimate main and interaction effects efficiently in industrial settings (fontana2023design; zubair2024coupling). Our work adopts this ANOVA/DoE perspective but targets pairwise ANOVA interaction maps of a fixed neural backbone and benchmarks interaction explanation methods, rather than proposing another globally additive predictive model.
@@
A growing line of work studies how sensitive explanations are to perturbations of inputs, parameters, or training conditions, and how consistent they remain across independently trained models. (agarwal2022rethinking) formalize stability metrics for attribution methods and show that popular saliency techniques can be highly unstable, while (xue2023stability) obtain formal stability guarantees by smoothing the predictor. Algorithmic approaches include training objectives that directly regularize explanation stability (chen2024training) and VAE-based schemes that generate more stable LIME-style surrogates (xiang2023stable). Beyond single models, (watson2022agree) show that identical architectures with different random seeds can yield markedly different explanations, (wu2024improving) propose an explanation-consistency score between original and adversarial images, and (rai8evaluating) quantify consistency of LIME, SHAP, and PDP explanations under varying class-imbalance regimes. These works primarily target local attributions or saliency maps and study stability under perturbations or retraining, whereas we focus on the consistency of pairwise interaction rankings and ANOVA interaction maps across independently trained backbones, measured via cross-backbone ranking and surface-alignment metrics.
@@
Several recent works explicitly evaluate how well interaction explanations recover known or controlled dependency structure. Examples include grammar-derived dependency relations with rank-based metrics in formal-language classifiers (jumelet2023feature), Rashomon-set analyses with interaction scores and Halo plots (li2023exploring), error-controlled discovery procedures with FDR guarantees on synthetic and semi-synthetic tasks (chen2025error), and stability-aware interaction evaluation on industrial intrusion-detection benchmarks (bouke2024application). XAI-Units and SPEX provide unit-test-style interaction benchmarks and LLM-scale setups with synthetic triggers and downstream QA tasks (lee2025xai; kang2025spex), while complementary work in NLP proposes unified evaluation frameworks for feature explanations with multiple faithfulness and robustness metrics (sun2025evaluating). In contrast, we develop a targeted evaluation protocol and metrics for global ANOVA interaction maps in regression backbones.
@@ 3.1 Problem setup and notation
We consider a trained regression backbone $f:\mathcal{X}\subset\mathbb{R}^{d}\to\mathbb{R}$, such as a multi-layer perceptron on tabular inputs or a small MLP on latent representations from a vision or language model. Let $X=(X_{1},\dots,X_{d})\sim P_{X}$ denote the input random vector and $\{x_{i}\}_{i=1}^{n}$ a sample drawn from $P_{X}$. Our goal is to construct, for each feature pair $(j,k)$, a low-dimensional ANOVA interaction map that summarizes how $f$ deviates from additivity in the $(X_{j},X_{k})$-plane, together with a scalar interaction strength that can be used for ranking and evaluation. Throughout this section, we treat $f$ and $P_{X}$ as fixed and focus on how to obtain these objects from black-box evaluations of $f$ on a discretized grid.
3.2 Target: ANOVA interaction map of a backbone
Given a fixed backbone $f$ and input distribution $P_{X}$, an ANOVA-style decomposition expresses the response as
@@ 3.2 Target: ANOVA interaction map of a backbone
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\bigl[f(X)\mid X_{j}=x_{j},X_{k}=x_{k}\bigr]$
$\quad-m_{j}(x_{j})-m_{k}(x_{k})-\mu.$
is the pairwise interaction component, with $r(x)$ collecting higher-order and residual terms. We refer to $g_{jk}$ as the ANOVA interaction map for the pair $(j,k)$, and to its $L^{2}$ norm
@@
as the corresponding target ANOVA interaction strengths. In practice, directly estimating the conditional expectations in these definitions is infeasible in moderate and high dimensions. Our method therefore fits a discrete factorial surrogate on a grid over $\mathcal{X}$ and recovers approximate ANOVA interaction maps and strengths from this surrogate.
@@ 3.3 Discretized factorial surrogate
We first discretize each feature into $L$ bins, inducing a regular grid over $\mathcal{X}$. For feature $j$, let $\mathcal{B}_{j}=\{1,\dots,L\}$ index bins, and let $q_{j}(x_{j})\in\mathcal{B}_{j}$ denote the bin assignment for value $x_{j}$ (e.g., via empirical quantiles or uniform spacing over the observed range). A grid point is thus identified by a level vector $\ell=(\ell_{1},\dots,\ell_{d})$ with $\ell_{j}\in\mathcal{B}_{j}$, and we write $f(\ell)$ as shorthand for the backbone prediction at a representative input from that cell.
On this grid, we construct an ANOVA-style linear surrogate
@@
where $\hat{m}_{j}:\mathcal{B}_{j}\to\mathbb{R}$ and $\hat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ are discrete main-effect and interaction tables. In matrix form, this corresponds to a least-squares fit of a factorial design matrix $\Phi\in\mathbb{R}^{n\times p}$, whose columns encode cell indicators and contrast-coded factors, to the backbone responses $\{f(x_{i})\}$. Our framework, ORACLE, solves
@@
where $\phi(q(x_{i}))$ denotes the ANOVA basis evaluated at the discretized levels $q(x_{i})=(q_{1}(x_{i1}),\dots,q_{d}(x_{id}))$. The coefficients $\hat{\beta}$ can then be reshaped into the tables $\hat{\mu}$, $\{\hat{m}_{j}\}$, and $\{\hat{g}_{jk}\}$.
3.4 Residual centering and $\mu$-rebalancing
A naive least-squares fit in the ANOVA basis does not necessarily enforce the ANOVA identifiability constraints, such as zero-mean main effects and zero row/column sums for interaction tables. To obtain a well-defined decomposition, we apply a simple post-processing step that recenters the estimated tables and rebalances the overall mean.
For each feature $j$, we subtract the empirical mean of the main-effect table to enforce
@@ 3.4 Residual centering and $\mu$-rebalancing
For each pair $(j,k)$, we adjust the interaction table $\hat{g}_{jk}$ so that its row and column averages vanish,
$\sum_{\ell_{j}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0,\quad\sum_{\ell_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0\quad\text{for all }\ell_{j},\ell_{k},$
while preserving the overall mean on the grid. Finally, any constant offset induced by these centering operations is absorbed into $\hat{\mu}$, a step we refer to as $\mu$-rebalancing. After this procedure, the discrete surrogate satisfies
$\hat{f}_{L}(\ell)\;=\;\hat{\mu}+\sum_{j=1}^{d}\hat{m}_{j}(\ell_{j})+\sum_{1\leq j<k\leq d}\hat{g}_{jk}(\ell_{j},\ell_{k})$
with the usual ANOVA orthogonality-style constraints at the level of the discretized grid. This makes the main and interaction tables interpretable as discrete approximations of the corresponding ANOVA components of $f$ under $P_{X}$. We summarize the centering and $\mu$-rebalancing procedure in Algorithm 1 in Appendix B.
3.5 Extracting ANOVA interaction maps and strengths
The post-processed interaction tables $\hat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ form discrete ANOVA interaction maps for each feature pair $(j,k)$. These maps can be visualized directly as heatmaps over the $(X_{j},X_{k})$ grid, revealing regions where the backbone exhibits strong non-additivity. To summarize interaction strength in a single scalar, we compute the root-mean-square (RMS) over the grid,

### v3

@@ Abstract
We introduce ORACLE, a framework for explaining neural networks on tabular data and scientific factorial designs. ORACLE summarizes a trained network’s prediction surface with main effects and pairwise interactions by treating the network as a black-box response, discretizing the inputs onto a grid, and fitting an orthogonal factorial (ANOVA-style) surrogate—the $L^{2}$ orthogonal projection of the model response onto a finite-dimensional factorial subspace. A simple centering and $\mu$-rebalancing step then expresses this surrogate as main- and interaction-effect tables that remain faithful to the original model in the $L^{2}$ sense. The resulting grid-based interaction maps are easy to visualize, comparable across backbones, and directly aligned with classical design-of-experiments practice. On synthetic factorial benchmarks and low- to medium-dimensional tabular regression tasks, ORACLE more accurately recovers ground-truth interaction structure and hotspots than Monte Carlo SHAP-family interaction methods, as measured by ranking, localization, and cross-backbone stability. In latent image and text settings, ORACLE clarifies its scope: grid-based factorial surrogates are most effective when features admit an interpretable factorial structure, making ORACLE particularly well-suited to scientific and engineering workflows that require stable, DoE-style interaction summaries.
@@ 1 Introduction
Modern neural networks often rely on rich interactions between features rather than purely additive structure, which makes them both powerful and difficult to interpret. Beyond marginal feature importance, practitioners increasingly need tools that describe how features act jointly, how such effects vary across the input space, and whether these patterns remain stable when the backbone model changes. This need is particularly acute in scientific and engineering tabular settings, where inputs represent physical or design variables and stakeholders value stable, low-order interaction summaries that connect naturally to classical design-of-experiments (DoE) practice. In this regime, functional ANOVA provides a natural formalism: it decomposes a response into main effects and interaction effects and, in classical DoE, factorial designs enable efficient estimation and visualization of low-order interactions under orthogonality and simple diagnostics.
In the neural-network setting, most interaction explanations have grown out of the SHAP literature rather than DoE, relying on Monte Carlo sampling over feature subsets and often lacking an explicit reference interaction surface. SHAP interaction indices and related methods can be computationally expensive, sensitive to sampling choices, and difficult to validate in terms of interaction detection, hotspot localization, and transfer across backbones. We revisit the DoE perspective and treat a trained neural network as a black-box response surface on which we fit a discrete ANOVA surrogate (i.e., an orthogonal factorial-effect expansion on a discretized input grid). Our framework, ORACLE (Orthogonal and Residual ANOVA for Concordant $L^{2}$-Effect Explanations), learns this orthogonal surrogate and applies recentering and $\mu$-rebalancing to obtain main-effect and pairwise interaction-effect tables that remain faithful to the backbone in an $L^{2}$ sense. The resulting low-dimensional interaction maps are directly visualizable, comparable across backbones, and aligned with DoE-style reasoning for tabular and scientific-design applications where features admit interpretable factorial structure and stakeholders value stable, low-order interaction summaries.
@@
    Formulate a discrete ANOVA surrogate for neural predictors that yields orthogonal main-effect and pairwise interaction-effect tables with explicit $L^{2}$ projection semantics and associated interaction maps.
@@
    Define target interaction strengths and oracle interaction maps via functional-ANOVA residuals on two-dimensional grids, providing a model-centric reference surface for evaluating interaction explanations.
@@
    Empirically compare ORACLE to SHAP-family interaction methods on synthetic factorial and real tabular regression benchmarks, highlighting regimes in which grid-based ANOVA surrogates yield more accurate and stable interaction explanations for low- to medium-dimensional tabular and scientific-design data.
@@ 2 Related Work
Recent work on global interaction explanations has largely built on the Shapley framework. Faith-Shap introduces an axiomatic Shapley interaction index as a canonical notion of global interaction strength (tsai2023faith), while SHAP-IQ, SVARM-IQ, and KernelSHAP-IQ provide Monte Carlo and kernel-based schemes to efficiently approximate any-order Shapley interactions in practice (fumagalli2023shap; kolpaczki2024svarm; fumagalli2024kernelshap). These algorithms and variants are consolidated and benchmarked in the shapiq library, which compares Shapley-based interaction estimators across models and datasets (muschalik2024shapiq). Complementary work analyzes how global feature effects decompose into main and interaction components (herbinger2024decomposing) and uses Shapley interactions to probe structured phenomena such as syntax and speech representations in large models (singhvi2024knowing). However, these approaches mostly return scalar interaction indices rather than explicit pairwise interaction maps, and are rarely evaluated against an interaction oracle or for stability across backbones.
ANOVA and design-of-experiments perspectives.
Functional ANOVA has re-emerged as a central tool for interpretable machine learning, motivating intrinsically interpretable models that decompose predictions into main effects and low-order interactions, as well as neural architectures that directly parameterize functional-ANOVA components or screen interactions before fitting a factorial surrogate, enabling higher-order terms with statistical guarantees (hu2025interpretable; kohler2025achieving; choi2025meta; park2025tensor). In engineering design and global sensitivity analysis, interpretable neural networks are combined with functional-ANOVA decompositions and grouped sensitivity indices to study complex physical interactions from observational data (dolar2024data). Complementary work in design-of-experiments couples factorial and fractional-factorial designs with modern machine learning to estimate main and interaction effects efficiently in industrial settings (fontana2023design; zubair2024coupling). Our work adopts this ANOVA/DoE perspective but targets pairwise interaction maps of a fixed neural backbone and benchmarks interaction explanation methods, rather than proposing another globally additive predictive model.
@@
A growing line of work studies how sensitive explanations are to perturbations of inputs, parameters, or training conditions, and how consistent they remain across independently trained models. (agarwal2022rethinking) formalize stability metrics for attribution methods and show that popular saliency techniques can be highly unstable, while (xue2023stability) obtain formal stability guarantees by smoothing the predictor. Algorithmic approaches include training objectives that directly regularize explanation stability (chen2024training) and VAE-based schemes that generate more stable LIME-style surrogates (xiang2023stable). Beyond single models, (watson2022agree) show that identical architectures with different random seeds can yield markedly different explanations, (wu2024improving) propose an explanation-consistency score between original and adversarial images, and (rai8evaluating) quantify consistency of LIME, SHAP, and PDP explanations under varying class-imbalance regimes. These works primarily target local attributions or saliency maps and study stability under perturbations or retraining, whereas we focus on the consistency of pairwise interaction rankings and interaction maps across independently trained backbones, measured via cross-backbone ranking and surface-alignment metrics.
@@
Several recent works explicitly evaluate how well interaction explanations recover known or controlled dependency structure. Examples include grammar-derived dependency relations with rank-based metrics in formal-language classifiers (jumelet2023feature), Rashomon-set analyses with interaction scores and Halo plots (li2023exploring), error-controlled discovery procedures with FDR guarantees on synthetic and semi-synthetic tasks (chen2025error), and stability-aware interaction evaluation on industrial intrusion-detection benchmarks (bouke2024application). XAI-Units and SPEX provide unit-test-style interaction benchmarks and LLM-scale setups with synthetic triggers and downstream QA tasks (lee2025xai; kang2025spex), while complementary work in NLP proposes unified evaluation frameworks for feature explanations with multiple faithfulness and robustness metrics (sun2025evaluating). In contrast, we develop a targeted evaluation protocol and metrics for global pairwise interaction maps in regression backbones.
@@ 3.1 Problem setup and notation
We consider a trained regression backbone $f:\mathcal{X}\subset\mathbb{R}^{d}\to\mathbb{R}$, such as a multi-layer perceptron on tabular inputs or a small MLP on latent representations from a vision or language model. Let $X=(X_{1},\dots,X_{d})\sim P_{X}$ denote the input random vector and $\{x_{i}\}_{i=1}^{n}$ a sample drawn from $P_{X}$. Our goal is to construct, for each feature pair $(j,k)$, a low-dimensional interaction map that summarizes how $f$ deviates from additivity on the $(X_{j},X_{k})$-plane, together with a scalar interaction strength used for ranking and evaluation. Throughout this section, we treat $f$ and $P_{X}$ as fixed and focus on obtaining these objects from black-box evaluations of $f$ on a discretized grid.
3.2 Target: interaction map of a backbone
Given a fixed backbone $f$ and input distribution $P_{X}$, the (functional) ANOVA decomposition expresses the response as
@@ 3.2 Target: interaction map of a backbone
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\bigl[f(X)\mid X_{j}=x_{j},X_{k}=x_{k}\bigr]-m_{j}(x_{j})-m_{k}(x_{k})-\mu$
is the pairwise interaction component, with $r(x)$ collecting higher-order and residual terms. We refer to $g_{jk}$ as the interaction map for the pair $(j,k)$, and to its $L^{2}$ norm
@@
as the corresponding target interaction strength. In practice, directly estimating the conditional expectations in these definitions is infeasible in moderate and high dimensions. ORACLE therefore fits a discrete factorial surrogate on a grid over $\mathcal{X}$ and recovers approximate interaction maps and strengths from this surrogate.
@@ 3.3 Discretized factorial surrogate
We discretize each feature into $L$ bins, inducing a regular grid over $\mathcal{X}$. For feature $j$, let $\mathcal{B}_{j}=\{1,\dots,L\}$ index bins, and let $q_{j}(x_{j})\in\mathcal{B}_{j}$ denote the bin assignment for value $x_{j}$ (e.g., via empirical quantiles or uniform spacing over the observed range). A grid point is identified by a level vector $\ell=(\ell_{1},\dots,\ell_{d})$ with $\ell_{j}\in\mathcal{B}_{j}$, and we write $f(\ell)$ as shorthand for the backbone prediction at a representative input from that cell.
On this grid, we fit a linear factorial surrogate
@@
where $\hat{m}_{j}:\mathcal{B}_{j}\to\mathbb{R}$ and $\hat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ are discrete main-effect and interaction effect tables. In matrix form, this corresponds to a least-squares fit of a factorial design matrix $\Phi\in\mathbb{R}^{n\times p}$, whose columns encode cell indicators and contrast-coded factors, to the backbone responses $\{f(x_{i})\}$. Our framework, ORACLE, solves
@@
where $\phi(q(x_{i}))$ denotes the basis evaluated at the discretized levels $q(x_{i})=(q_{1}(x_{i1}),\dots,q_{d}(x_{id}))$. The coefficients $\hat{\beta}$ can then be reshaped into the tables $\hat{\mu}$, $\{\hat{m}_{j}\}$, and $\{\hat{g}_{jk}\}$.
3.4 Recentering and $\mu$-rebalancing (identifiability)
A naive least-squares fit does not necessarily satisfy the standard identifiability constraints for factorial effect tables, such as zero-mean main effects and zero row/column sums for interaction tables. To obtain a well-defined decomposition, we apply a lightweight post-processing step that recenters the estimated tables and rebalances the overall mean.
For each feature $j$, we subtract the average of the main-effect table to enforce
@@ 3.4 Recentering and $\mu$-rebalancing (identifiability)
For each pair $(j,k)$, we adjust the interaction table $\hat{g}_{jk}$ so that its row and column sums vanish,
$\sum_{\ell_{j}\in\mathcal{B}_{j}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0,\quad\sum_{\ell_{k}\in\mathcal{B}_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0\quad\text{for all }\ell_{k}\in\mathcal{B}_{k},\ \ell_{j}\in\mathcal{B}_{j},$
while preserving the overall mean on the grid. Finally, any constant offset induced by these centering operations is absorbed into $\hat{\mu}$ ($\mu$-rebalancing”). After this procedure, the discrete surrogate satisfies
with the usual sum-to-zero constraints at the discretized-grid level. This makes the effect tables interpretable as discrete approximations of the corresponding functional-ANOVA components of $f$ under $P_{X}$. We summarize the recentering and $\mu$-rebalancing procedure in Algorithm 1 (Appendix B).
3.5 Extracting interaction maps and strengths
The post-processed interaction tables $\hat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ form discrete interaction maps for each feature pair $(j,k)$. These maps can be visualized directly as heatmaps over the $(X_{j},X_{k})$ grid, revealing regions where the backbone exhibits strong non-additivity. To summarize interaction strength in a single scalar, we compute the root-mean-square (RMS) over the grid,